You are the pricing manager for a new longboard manufacturing company. You are getting ready to launch your first longboard into the market. Your cost to manufacture the longboard is $80. You are tasked with pricing the longboard to assure a gross profit margin (margin) of 40%. Please show your calculation and proposed price for the product. with pricing the longboard to assure a gross profit margin (margin) of 40%. Please show your calculation and proposed price for the product.
Answer #1

Gross Profit Margin = ((Revenue or selling price – Cost of Goods Sold or Manufacturing cost)/Cost of Goods Sold)*100%

Given:

Manufacturing Cost = $80

Gross Profit Margin = 40% = 0.4

Selling price we have to find = x (let)

((x – 80)/80)*100 = 40

x – 80 = 40*80/100 = 32

x = 32 + 80 = 112

Selling Price = x = $112

Hence,

Pricing the long board at $112 will give a gross profit margin of 40%.

Proposed price for the product is $112

********

IF YOU HAVE ANY DOUBT, KINDLY COMMENT. I WOULD LOVE TO HELP YOU!!!

IF YOU LIKE THE ANSWER, PLEASE GIVE AN UP-VOTE OR THUMB UP. THIS WILL ENCOURAGE ME TO ANSWER MORE!!